OpenStudy (anonymous):

what does a number in front of a logarithm do? (example: 2 log(100) )

OpenStudy (anonymous):

It means that the value is to a power that is equal to that coefficient. So... log(100)^2

OpenStudy (anonymous):

rogster is absolutely right: the law of logarithm says that : \[n log(m) = log(m)^n\] where n is a number

OpenStudy (amistre64):

it multiplies the value of the log ....

OpenStudy (amistre64):

2 log(100) = 2 * 2 = 4